OpenStudy (anonymous):

How many grams of methane gas (CH4) are in a 15.8 liter sample at 1.2 atmospheres and 27°C? Show all work used to solve this problem.

OpenStudy (aaronq):

first find the moles (n) of gas present, use PV=nRT you need to convert the temperature from Celsius to Kelvin.

OpenStudy (anonymous):

Is that when you add 274 to 27?

OpenStudy (aaronq):

yeah 273 + 27 R in these units is 0.08206

OpenStudy (aaronq):

After you've found the moles you wanna convert them to grams \(\sf moles=\dfrac{mass}{molar ~mass}\)

OpenStudy (anonymous):

how do you finds the moles?

OpenStudy (somy):

PV= nRT P= pressure V= volume n= mole R= constant = 0.08206 T= temperature

OpenStudy (somy):

put the values into the formula and get the mole

OpenStudy (anonymous):

what would p be?

OpenStudy (anonymous):

I am so confused I have literally been staring at this problem and another one for and hour trying to figure it out.

OpenStudy (aaronq):

P is pressure, it varies in units, in this problem it's given as atmospheres (ATM)

OpenStudy (anonymous):

oh okay. so it would be set up like 1.2 atm x 15.8 liter = n0.08206 x 300

OpenStudy (anonymous):

would the answer be 0.77?

OpenStudy (somy):

yes

OpenStudy (anonymous):

Thank you!

OpenStudy (somy):

oh u r not finished tho lol :)

OpenStudy (anonymous):

Oh aha then what do I do now?

OpenStudy (anonymous):

would i use stoich and convert the mols to grams?

OpenStudy (anonymous):

would the final answer be 0.048

OpenStudy (somy):

nope

OpenStudy (somy):

look mole= mass/ Molecular mass so mass= mole* Molecular mass you got mole = 0.77 molecular mass of CH4= 12+4= 16 thus mass= 0.77* 16=

OpenStudy (somy):

the answer will be your mass